# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ Acknowledgements
[…]
Appendix A Proof for Section 2
@@
For $d=1$, the matrix representation of $K_{s}\in\mathbb{R}^{|\mathsf{Chi}(s)|\times m}$ is $K_{s}=(I-\mathbf{1}(\frac{|\mathsf{Des}(j)|}{|\mathsf{Des}(s)|})_{j\in\mathsf{Chi}(s)}^{T})P_{s}$ where $P_{s}\in\mathbb{R}^{|\mathsf{Chi}(s)|\times m}$, and $P_{s}(i,j)=|\mathsf{Des}(i)|^{-1}$ if $j\in\mathsf{Des}(i)$ and $P_{s}(i,j)=0$ otherwise. Notice that
@@
If $s$ is not in the subtree rooted at $s^{\prime}$ and $s^{\prime}$ is not in the subtree rooted at $s$, then $\mathsf{Des}(s)\cap\mathsf{Des}(s^{\prime})=\emptyset$, and so $P_{s}P_{s^{\prime}}^{T}=\mathbf{0}$. If $s$ is in the subtree rooted at $s^{\prime}$, then $s$ is in the subtree rooted at some $\widehat{s}\in\mathsf{Chi}(s^{\prime})$. In particular, $(P_{s}P_{s^{\prime}}^{T})(i,j)=0$ for all $j\neq\widehat{s}$ and $(P_{s}P_{s^{\prime}}^{T})(i,\widehat{s})=|\mathsf{Des}(\widehat{s})|^{-1}$, and thus $(I-\mathbf{1}(\frac{|\mathsf{Des}(j)|}{|\mathsf{Des}(s)|})_{j\in\mathsf{Chi}(s)}^{T})P_{s}P_{s^{\prime}}^{T}=\mathbf{0}$. Similarly for the case when $s^{\prime}$ is in the subtree rooted at $s$. The case when $d>1$ follows by applying the above argument coordinate-wise.
For the second claim, consider the case $d=1$, denoting
@@
where the norm in the denominator in the definition of $v$ is the $l_{2}$ norm. Then when $d=1$, we have (applying Theorem 6 in [Meyer1973Generalized])
@@
The above argument can be applied coordinate-wise, and so extend to $d\geq 1$. ∎
Appendix B Proofs for Section 3.2
Lemma B.1 (generalized lemma 5 in [lan_communication-efficient_2020]).
$u^{*}\in\operatorname*{argmin}_{u\in U}q(u)+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,x_{i}).$
Then for any $u\in U$, we have
$q(u^{*})+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u^{*},x_{i})\leq q(u)+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,x_{i})-\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,u^{*}).$
Proof of Lemma B.1.
$\langle q^{\prime}(u^{*})+\sum_{i\in\mathcal{I}}\eta_{i}\nabla D_{w}(u^{*},x_{i}),u-u^{*}\rangle\geq 0,\quad\forall u\in U.$
By definition, we have for each $i\in\mathcal{I}$ that
$D_{w}(u,x_{i})-D_{w}(u^{*},x_{i})-D_{w}(u,u^{*})=\langle\nabla w(x_{i})-\nabla w(u^{*}),u-u^{*}\rangle=-\langle\nabla D_{w}(u^{*},x_{i}),u-u^{*}\rangle$
Thus, we have for any $u\in U$,
$\quad q(u)+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,x_{i})$
$\geq q(u^{*})+\langle q^{\prime}(u^{*}),u-u^{*}\rangle+\sum_{i\in\mathcal{I}}\eta_{i}\left(D_{w}(u^{*},x_{i})+D_{w}(u,u^{*})-\langle\nabla D_{w}(u^{*},x_{i}),u-u^{*}\rangle\right)$
$\geq q(u^{*})+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u^{*},x_{i})+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,u^{*}).$
Lemma B.2.
then for $t\geq 1$ and $u\in U$
$(\sum_{t=1}^{T}\lambda_{t})\cdot(\Phi(\widehat{u}^{T})-\Phi(u))\leq(\eta\beta_{1}-\mu/C)\lambda_{1}D_{w}(u,u^{0})-\eta(1+\beta_{T})\lambda_{T}D_{w}(u,u^{T})+\sum_{t=1}^{T}\frac{M^{2}\lambda_{t}}{2\eta\beta_{t}}.$
Proof of Lemma B.2.
Applying Lemma B.1, and using $\sum_{i\in\mathcal{I}}\eta_{i}=\eta$, we have
$\quad\langle v+\phi^{\prime}(u^{t-1}),u^{t}-u\rangle+\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u^{t},x_{i})-\sum_{i\in\mathcal{I}}\eta_{i}D_{w}(u,x_{i})$
$\leq\eta\beta_{t}D_{w}(u,u^{t-1})-\eta\beta_{t}D_{w}(u^{t},u^{t-1})-(1+\beta_{t})\eta D_{w}(u,u^{t})$
The rest follows a similar argument as in the proof of Proposition 2 [lan_communication-efficient_2020].
Appendix C Proofs for Section 3.3
Proof of Lemma 3.1.
For convenience, denote $\eta_{k}=\sum_{s=1}^{S}\eta_{k,s}$ and $\overline{r}=\sum_{s=1}^{S}r_{s}\rho_{s}$.
Primal update properties. By Corollary 3.1, for any $v\in V$ and $x_{v}\in\mathcal{X}$
$\quad\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},\widehat{x}_{v}^{k}-x_{v}\rangle+f_{v}(\widehat{x}_{v}^{k})-f_{v}(x_{v})$
$\leq\frac{2\eta_{k}}{T_{k}(T_{k}+3)}D_{w^{x}}(x_{v},x_{v}^{k-1})+\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(x_{v},x_{v}^{k-r_{s}})$
$\quad-\frac{(T_{k}+1)(T_{k}+2)}{T_{k}(T_{k}+3)}\eta_{k}D_{w^{x}}(x_{v},x_{v}^{k})-\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(\widehat{x}_{v}^{k},x_{v}^{k-r_{s}})+\frac{4M^{2}}{\eta_{k}(T_{k}+3)}.$
Summing over $k$, and defining $\eta_{k,s}=0$ for all $k<0$ and $k\geq N+1$, we have
$\quad\sum_{k=0}^{N}\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},\widehat{x}_{v}^{k}-x_{v}\rangle+f_{v}(\widehat{x}_{v}^{k})-f_{v}(x_{v})$
$\leq\sum_{k=0-\max\{r_{s}\}}^{N}\left(\frac{2\eta_{k+1}}{T_{k+1}(T_{k+1}+3)}+\sum_{s=1}^{S}\eta_{k+r_{s},s}\right)D_{w^{x}}(x_{v},x_{v}^{k})$
$\quad-\sum_{k=0}^{N}\frac{(T_{k}+1)(T_{k}+2)}{T_{k}(T_{k}+3)}\eta_{k}D_{w^{x}}(x_{v},x_{v}^{k})$
$\quad-\sum_{k=0}^{N}\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(\widehat{x}_{v}^{k},x_{v}^{k-r_{s}})+\sum_{k=0}^{N}\frac{4M^{2}}{\eta_{k}(T_{k}+3)}.$
Recall that $x_{v}^{k}=x_{v}^{init}$ for all $k<0$, and so with $T_{k}=T$ for all $k$ and $\eta_{k,s}=\eta\rho_{s}$ for $k=0,1,\ldots,N$,
$\sum_{k=0-\max\{r_{s}\}}^{-1}\left(\frac{2\eta_{k+1}}{T(T+3)}+\sum_{s=1}^{S}\eta_{k+r_{s},s}\right)\leq\eta(\frac{2}{T(T+3)}+\sum_{s=1}^{S}r_{s}\rho_{s})=\eta(\frac{2}{T(T+3)}+\overline{r}).$
For $k=0,1,\ldots,N-1$
$\frac{2\eta_{k+1}}{T_{k+1}(T_{k+1}+3)}+\sum_{s=1}^{S}\eta_{k+r_{s},s}\leq\eta(\frac{2}{T(T+3)}+\sum_{s=1}^{S}\rho_{s})=\frac{\eta\cdot(T+1)(T+2)}{T(T+3)}.$
Thus we have
$\quad\sum_{k=0}^{N}\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},\widehat{x}_{v}^{k}-x_{v}\rangle+f_{v}(\widehat{x}_{v}^{k})-f_{v}(x_{v})$
$\leq\eta(\frac{2}{T(T+3)}+\overline{r})D_{w^{x}}(x_{v},x_{v}^{init})-\frac{(T+1)(T+2)}{T(T+3)}\eta D_{w^{x}}(x_{v},x_{v}^{N})$
$\quad-\eta\sum_{k=0}^{N}\sum_{s=1}^{S}\rho_{s}D_{w^{x}}(\widehat{x}_{v}^{k},x_{v}^{k-r_{s}})+\frac{4M^{2}(N+1)}{\eta(T+3)}$ (25)
Dual update properties. In addition, by the updating rule for the dual, we have by Proposition 2 in [lan_communication-efficient_2020] for any $y_{s}\in\mathbb{R}^{n_{s}}$,
$\quad\langle-\frac{1}{r_{s}}\sum_{v\in V}K_{s,v}\widetilde{x}_{s,v}^{i_{s}},y_{s}^{i_{s}}-y_{s}\rangle+R^{*}_{s}(y_{s}^{i_{s}})-R^{*}_{s}(y_{s})$
$\leq\tau_{s,i_{s}}\left(D_{w^{y}_{s}}(y_{s},y_{s}^{i_{s}-1})-D_{w^{y}_{s}}(y_{s},y_{s}^{i_{s}})-D_{w^{y}_{s}}(y^{i_{s}}_{s},y_{s}^{i_{s}-1})\right).$
$\quad\sum_{i_{s}=0}^{N_{s}-1}\langle-\frac{1}{r_{s}}\sum_{v\in V}K_{s,v}\widetilde{x}_{s,v}^{i_{s}},y_{s}^{i_{s}}-y_{s}\rangle+R^{*}_{s}(y_{s}^{i_{s}})-R^{*}_{s}(y_{s})$
$\leq\tau_{s}\left(D_{w^{y}_{s}}(y_{s},y_{s}^{init})-D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})\right)-\tau_{s}\cdot\sum_{i_{s}=0}^{N_{s}-1}D_{w^{y}_{s}}(y^{i_{s}}_{s},y_{s}^{i_{s}-1}).$ (26)
Gap properties. Notice that for each $s\in[S]$ and $v\in V$, we have
$\quad\sum_{k=0}^{N}\left\{\langle\widehat{x}_{v}^{k},K^{*}_{s,v}y_{s}\rangle-\langle x_{v},K^{*}_{s,v}\overline{y}_{s}^{k}\rangle\right\}$
$=\sum_{i_{s}=0}^{N_{s}-1}\left\{\sum_{i=0}^{r_{s}-1}\langle\widehat{x}_{v}^{r_{s}i_{s}+i},K^{*}_{s,v}y_{s}\rangle-r_{s}\langle x_{v},K^{*}_{s,v}y_{s}^{i_{s}}\rangle\right\}$
$=\sum_{i_{s}=0}^{N_{s}-1}\langle\sum_{i=0}^{r_{s}-1}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle+\sum_{k=0}^{N}\langle\widehat{x}_{v}^{k}-x_{v},K^{*}_{s,v}\overline{y}_{s}^{k}\rangle+\sum_{i_{s}=0}^{N_{s}-1}\langle K_{s,v}\widetilde{x}_{s,v}^{i_{s}},y_{s}-y_{s}^{i_{s}}\rangle.$ (27)
Recall that for $i_{s}=0,1,\ldots,N_{s}-1$,
$\widetilde{x}_{s,v}^{i_{s}}=\alpha(\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}\widehat{x}_{v}^{k^{\prime}}-\sum_{k^{\prime}=r_{s}i_{s}-2r_{s}}^{r_{s}i_{s}-r_{s}-1}{x}_{v}^{k^{\prime}})+\sum_{k^{\prime}=r_{s}i_{s}-r_{s}}^{r_{s}i_{s}-1}{x}_{v}^{k^{\prime}}.$
We first bound the first term in (C). Notice that for $i_{s}=0,1,\ldots,N_{s}-1$, we have
$\quad\langle\sum_{i=0}^{r_{s}-1}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
$=\langle\sum_{i=0}^{r_{s}-1}(\widehat{x}_{v}^{r_{s}i_{s}+i}-x_{v}^{r_{s}(i_{s}-1)+i})-\alpha\sum_{i=0}^{r_{s}-1}(\widehat{x}_{v}^{r_{s}(i_{s}-1)+i}-x_{v}^{r_{s}(i_{s}-2)+i}),K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
$=\langle\sum_{i=0}^{r_{s}-1}(\widehat{x}_{v}^{r_{s}i_{s}+i}-x_{v}^{r_{s}(i_{s}-1)+i}),K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
$\quad-\alpha\langle\sum_{i=0}^{r_{s}-1}(\widehat{x}_{v}^{r_{s}(i_{s}-1)+i}-x_{v}^{r_{s}(i_{s}-2)+i}),K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}-1})\rangle$
$\quad+\alpha\langle\sum_{i=0}^{r_{s}-1}(\widehat{x}_{v}^{r_{s}(i_{s}-1)+i}-x_{v}^{r_{s}(i_{s}-2)+i}),K^{*}_{s,v}(y_{s}^{i_{s}}-y_{s}^{i_{s}-1})\rangle.$
Thus, with $\alpha=1$, and recall that for $i_{s}=0$, $i=0,\ldots,r_{s}-1$, $\widehat{x}_{v}^{r_{s}(i_{s}-1)+i}-x_{v}^{r_{s}(i_{s}-2)+i}=x_{v}^{init}-x_{v}^{init}=\mathbf{0}$, we have
$\quad\sum_{v\in V}\sum_{i_{s}=0}^{N_{s}-1}\langle\sum_{i=0}^{r_{s}-1}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
Thus, for any $\rho>0$, we have
$\quad\sum_{v\in V}\sum_{i_{s}=0}^{N_{s}-1}\langle\sum_{i=0}^{r_{s}-1}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
Bounding the gap. Thus, with (C), (C), and (C), we have the following upper bound on the gap
$\quad\sum_{k=0}^{N}Q(\widehat{X}^{k},\overline{Y}^{k};Z)$
$\leq\eta(\frac{2}{T(T+3)}+\overline{r})D_{w^{X}}(X,X^{init})-\frac{(T+1)(T+2)}{T(T+3)}\eta D_{w^{X}}(X,X^{N})$
$\quad-\eta\sum_{k=0}^{N}\sum_{s=1}^{S}\rho_{s}D_{w^{X}}(\widehat{X}^{k},X^{k-r_{s}})+\frac{4mM^{2}(N+1)}{\eta(T+3)}$
$\quad+\sum_{s=1}^{S}\tau_{s}r_{s}\left\{D_{w^{y}_{s}}(y_{s},y_{s}^{init})-D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})-\sum_{i_{s}=0}^{N_{s}-1}D_{w^{y}_{s}}(y^{i_{s}}_{s},y_{s}^{i_{s}-1})\right\}$
where we take $\rho=\eta\rho_{s}$ in (C). Thus, with $\frac{\widetilde{\kappa}^{2}_{s}}{\rho_{s}\tau_{s}}\leq\frac{\eta}{2}$ for all $s\in[S]$, we have
$\quad\sum_{k=0}^{N}Q(\widehat{X}^{k},\overline{Y}^{k};Z)$
$\leq\eta(\frac{2}{T(T+3)}+\overline{r})D_{w^{X}}(X,X^{init})-\frac{(T+1)(T+2)}{T(T+3)}\eta D_{w^{X}}(X,X^{N})$
$\quad+\sum_{s=1}^{S}\tau_{s}r_{s}\cdot\left\{\frac{3}{2}D_{w^{y}_{s}}(y_{s},y_{s}^{init})-\frac{1}{2}D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})\right\}+\frac{4mM^{2}(N+1)}{\eta(T+3)}$
In particular, taking $\tau_{s}=\frac{2\widetilde{\kappa}_{s}^{2}}{\rho_{s}\eta}$, and using the convexity, we get
$\quad(N+1)\cdot Q(Z^{N};Z)$
$\leq\eta\left\{(\frac{2}{T(T+3)}+\overline{r})D_{w^{X}}(X,X^{init})-\frac{(T+1)(T+2)}{T(T+3)}D_{w^{X}}(X,X^{N})\right\}$
$\quad+\frac{1}{\eta}\left\{\sum_{s=1}^{S}\frac{\widetilde{\kappa}_{s}^{2}r_{s}}{\rho_{s}}\left\{3D_{w^{y}_{s}}(y_{s},y_{s}^{init})-D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})\right\}+\frac{4mM^{2}(N+1)}{T+3}\right\}$
The result follows from noticing that for any $T\geq 1$, $\frac{2}{T(T+3)}\leq\frac{1}{2}\leq\frac{1}{2}(\sum_{s=1}^{S}r_{s}\rho_{s})$, and $\frac{(T+1)(T+2)}{T(T+3)}\geq 1$. ∎
Appendix D Proofs for Section 4
Proof of Lemma 4.1.
Primal update properties. By Corollary 3.1, for any $v\in V$ and $x_{v}\in\mathcal{X}$
$\quad\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},\widehat{x}_{v}^{k}-x_{v}\rangle+f_{v}(\widehat{x}_{v}^{k})-f_{v}(x_{v})$
$\leq\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(x_{v},x_{v}^{k-r_{s}})-\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(\widehat{x}_{v}^{k},x_{v}^{k-r_{s}})$
$\quad\quad-(\frac{\mu}{C}+\eta_{k})D_{w^{x}}(x_{v},x_{v}^{k})+\frac{2M^{2}}{\eta_{k}T_{k}(T_{k}+1)}\sum_{t=1}^{T_{k}}\frac{\lambda_{t}}{\beta_{t}^{k}}.$
Thus, taking a weighted sum of the above, we get
$\quad\sum_{k=0}^{N}\theta_{k}\left\{\langle\sum_{s=1}^{S}K_{s,v}^{*}\overline{y}_{s}^{k},\widehat{x}_{v}^{k}-x_{v}\rangle+f_{v}(\widehat{x}_{v}^{k})-f_{v}(x_{v})\right\}$
$\leq\frac{\mu(\overline{r^{3}}/\overline{r}+5(\overline{r^{2}}/\overline{r})^{2})}{2C}D_{w^{x}}(x_{v},x_{v}^{init})-\sum_{k=0}^{N}\theta_{k}\sum_{s=1}^{S}\eta_{k,s}D_{w^{x}}(\widehat{x}_{v}^{k},x_{v}^{k-r_{s}})+\frac{\mu\overline{r^{2}}}{C\overline{r}}D_{0},$
where the bounds on the sum of Bregman divergence terms follows from the following bounds on the coefficients of $D_{w^{x}}(x_{v},x_{v}^{k})$ term in the sum ($\theta_{k}=\eta_{k}=0$ for all $k\geq N+1$) for $k=0,1,\ldots,N$
and for $x_{v}^{init}$
The bound on the rest of the terms is since
and notice that
and since $N+1\geq r_{\max}:=\max_{s\in[S]}r_{s}$, we have
$\leq\frac{2M^{2}C}{\mu}\left\{\frac{10N^{2}}{T^{2}}+\frac{32N\overline{r}}{T}\right\}\leq\frac{\mu}{C}D_{0}(\overline{r^{2}}/\overline{r})^{2}$
Dual update properties. Similar to (C), we get
$\quad\sum_{i_{s}=0}^{N_{s}-1}\left\{\langle-\sum_{v\in V}K_{s,v}\widetilde{x}_{s,v}^{i_{s}},y_{s}^{i_{s}}-y_{s}\rangle+(\sum_{k^{\prime}=r_{s}i_{s}}^{r_{s}i_{s}+r_{s}-1}\theta_{k}^{\prime})(R^{*}_{s}(y_{s}^{i_{s}})-R^{*}_{s}(y_{s}))\right\}$
$\leq\sum_{i_{s}=0}^{N_{s}-1}\tau_{s,i_{s}}(\sum_{k^{\prime}=r_{s}i_{s}}^{r_{s}i_{s}+r_{s}-1}\theta_{k}^{\prime})\left\{D_{w^{y}_{s}}(y_{s},y_{s}^{i_{s}-1})-D_{w^{y}_{s}}(y_{s},y_{s}^{i_{s}})-D_{w^{y}_{s}}(y^{i_{s}}_{s},y_{s}^{i_{s}-1})\right\}$
$=\tau_{s}\left\{D_{w^{y}_{s}}(y_{s},y_{s}^{init})-D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})-\sum_{i_{s}=0}^{N_{s}-1}D_{w^{y}_{s}}(y^{i_{s}}_{s},y_{s}^{i_{s}-1})\right\}.$ (29)
Gap properties. Notice that for each $s\in[S]$ and $v\in V$, we have
$\quad\sum_{k=0}^{N}\theta_{k}\left\{\langle\widehat{x}_{v}^{k},K^{*}_{s,v}y_{s}\rangle-\langle x_{v},K^{*}_{s,v}\overline{y}_{s}^{k}\rangle\right\}$
$=\sum_{i_{s}=0}^{N_{s}-1}\langle\sum_{i=0}^{r_{s}-1}\theta_{r_{s}i_{s}+i}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
$\quad+\sum_{k=0}^{N}\theta_{k}\langle\widehat{x}_{v}^{k}-x_{v},K^{*}_{s,v}\overline{y}_{s}^{k}\rangle+\sum_{i_{s}=0}^{N_{s}-1}\langle K_{s,v}\widetilde{x}_{s,v}^{i_{s}},y_{s}-y_{s}^{i_{s}}\rangle.$ (30)
We first bound the first term in (C). With $\alpha_{s,i_{s}}=1$
$\quad\sum_{v\in V}\sum_{i_{s}=0}^{N_{s}-1}\langle\sum_{i=0}^{r_{s}-1}\theta_{r_{s}i_{s}+i}\widehat{x}_{v}^{r_{s}i_{s}+i}-\widetilde{x}_{s,v}^{i_{s}},K^{*}_{s,v}(y_{s}-y_{s}^{i_{s}})\rangle$
$\quad\sum_{k=0}^{N}\theta_{k}Q(\widehat{X}^{k},\overline{Y}^{k};Z)$
$\leq\frac{\mu(\overline{r^{3}}/\overline{r}+5(\overline{r^{2}}/\overline{r})^{2})}{2C}D_{w^{X}}(X,X^{init})-\sum_{k=0}^{N}\theta_{k}\sum_{s=1}^{S}\eta_{k,s}D_{w^{X}}(\widehat{X}^{k},X^{k-r_{s}})+\frac{m\mu(\overline{r^{2}}/\overline{r})^{2}}{C}D_{0}$
$\quad+\sum_{s=1}^{S}\tau_{s}\left\{D_{w^{y}_{s}}(y_{s},y_{s}^{init})-D_{w^{y}_{s}}(y_{s},y_{s}^{N_{s}-1})-\sum_{i_{s}=1}^{N_{s}-1}D_{w^{y}_{s}}(y^{i_{s}}_{s},y_{s}^{i_{s}-1})\right\}$
$\leq\frac{\mu(\overline{r^{3}}/\overline{r}+5(\overline{r^{2}}/\overline{r})^{2})}{2C}D_{w^{X}}(X,X^{init})+\frac{m\mu(\overline{r^{2}}/\overline{r})^{2}}{C}D_{0}+\sum_{s=1}^{S}\tau_{s}D_{w^{y}_{s}}(y_{s},y_{s}^{init})$
using
Since $\sum_{k=0}^{N}\theta_{k}\geq\frac{N(N+1)}{2}$, using convexity, we get
[…]
$\quad+\frac{m\mu(\overline{r^{2}}/\overline{r})^{2}}{C}D_{0}+\frac{4\overline{r}C}{\mu}\sum_{s=1}^{S}\frac{\widetilde{\kappa}_{s}^{2}r_{s}}{\rho_{s}}D_{w^{y}_{s}}(y_{s},y_{s}^{init})\biggr{\}}$

### v2

@@ Acknowledgements
Appendix A Generalized gradient sliding procedure
In Algorithm 3, we provide the generalized gradient sliding procedure. Then we provide the proofs of its properties.
Algorithm 3 Generalized gradient sliding procedure
$(u^{T},\widehat{u}^{T})=GS(\phi,U,D,T,(\eta_{i})_{i\in\mathcal{I}},v,(x_{i})_{i\in\mathcal{I}},x^{init})$, an approximate solution to
$\min_{u\in U}\Phi(u):=\langle v,u\rangle+\phi(u)+\sum_{i\in\mathcal{I}}\eta_{i}D(u,x_{i})$
for $t=1,\ldots,T$ do
$u^{t}=\operatorname*{argmin}_{u\in U}\langle v+\phi^{\prime}(u^{t-1}),u\rangle+\sum_{i\in\mathcal{I}}\eta_{i}D(u,x_{i})+\eta\beta_{t}D(u,u^{t-1})$
end for
$\widehat{u}^{T}=(\sum_{t=1}^{T}\lambda_{t})^{-1}\sum_{t=1}^{T}\lambda_{t}u^{t}.$
Lemma A.1 (generalized lemma 5 in [27]).
$u^{*}\in\operatorname*{argmin}_{u\in U}q(u)+\sum_{i\in\mathcal{I}}\eta_{i}D(u,x_{i}).$
Then for any $u\in U$, we have
$q(u^{*})+\sum_{i\in\mathcal{I}}\eta_{i}D(u^{*},x_{i})\leq q(u)+\sum_{i\in\mathcal{I}}\eta_{i}D(u,x_{i})-\sum_{i\in\mathcal{I}}\eta_{i}D(u,u^{*}).$
Proof of Lemma A.1.
$\langle q^{\prime}(u^{*})+\sum_{i\in\mathcal{I}}\eta_{i}\nabla D(u^{*},x_{i}),u-u^{*}\rangle\geq 0,\quad\forall u\in U.$
By definition, we have for each $i\in\mathcal{I}$ that
$D(u,x_{i})-D(u^{*},x_{i})-D(u,u^{*})=\langle\nabla w(x_{i})-\nabla w(u^{*}),u-u^{*}\rangle=-\langle\nabla D(u^{*},x_{i}),u-u^{*}\rangle$
Thus, we have for any $u\in U$,
$\quad q(u)+\sum_{i\in\mathcal{I}}\eta_{i}D(u,x_{i})$
$\geq q(u^{*})+\langle q^{\prime}(u^{*}),u-u^{*}\rangle+\sum_{i\in\mathcal{I}}\eta_{i}\left(D(u^{*},x_{i})+D_{w}(u,u^{*})-\langle\nabla D(u^{*},x_{i}),u-u^{*}\rangle\right)$
$\geq q(u^{*})+\sum_{i\in\mathcal{I}}\eta_{i}D(u^{*},x_{i})+\sum_{i\in\mathcal{I}}\eta_{i}D(u,u^{*}).$
Lemma A.2.
then for $t\geq 1$ and $u\in U$
$(\sum_{t=1}^{T}\lambda_{t})\cdot(\Phi(\widehat{u}^{T})-\Phi(u))\leq(\eta\beta_{1}-\mu/C)\lambda_{1}D(u,u^{0})-\eta(1+\beta_{T})\lambda_{T}D(u,u^{T})+\sum_{t=1}^{T}\frac{M^{2}\lambda_{t}}{2\eta\beta_{t}}.$
Proof of Lemma A.2.
Applying Lemma A.1, and using $\sum_{i\in\mathcal{I}}\eta_{i}=\eta$, we have
$\quad\langle v+\phi^{\prime}(u^{t-1}),u^{t}-u\rangle+\sum_{i\in\mathcal{I}}\eta_{i}D(u^{t},x_{i})-\sum_{i\in\mathcal{I}}\eta_{i}D(u,x_{i})$
$\leq\eta\beta_{t}D(u,u^{t-1})-\eta\beta_{t}D(u^{t},u^{t-1})-(1+\beta_{t})\eta D(u,u^{t})$
The rest follows a similar argument as in the proof of Proposition 2 [27].
As a corollary to Lemma A.2, we have the following performance guarantee.
Corollary A.1.
$\langle v,\widehat{u}^{T}-u\rangle+\phi(\widehat{u}^{T})-\phi(u)$ $\leq\frac{2\eta}{T(T+3)}D(u,x^{init})+\sum_{i\in\mathcal{I}}\eta_{i}D({u},x_{i})$
$\quad-\frac{(T+1)(T+2)}{T(T+3)}\eta D(u,u^{T})-\sum_{i\in\mathcal{I}}\eta_{i}D(\widehat{u}^{T},x_{i})+\frac{4M^{2}}{\eta(T+3)}.$
$\langle v,\widehat{u}^{T}-u\rangle+\phi(\widehat{u}^{T})-\phi(u)$ $\leq\sum_{i\in\mathcal{I}}\eta_{i}D({u},x_{i})-\sum_{i\in\mathcal{I}}\eta_{i}D(\widehat{u}^{T},x_{i})$
$\quad-(\frac{\mu}{C}+\eta)D(u,u^{T})+\frac{2M^{2}/\eta}{T(T+1)}\sum_{t=1}^{T}\frac{\lambda_{t}}{\beta_{t}},$
Appendix B Proof for Section 4.2
@@ Appendix A Generalized gradient sliding procedure
For $\overline{d}=1$, the matrix representation of $K_{s}\in\mathbb{R}^{|\mathsf{Chi}(s)|\times m}$ is $K_{s}=(I-\mathbf{1}(\frac{|\mathsf{Des}(j)|}{|\mathsf{Des}(s)|})_{j\in\mathsf{Chi}(s)}^{T})P_{s}$ where $P_{s}\in\mathbb{R}^{|\mathsf{Chi}(s)|\times m}$, and $P_{s}(i,j)=|\mathsf{Des}(i)|^{-1}$ if $j\in\mathsf{Des}(i)$ and $P_{s}(i,j)=0$ otherwise. Notice that
@@
If $s$ is not in the subtree rooted at $s^{\prime}$ and $s^{\prime}$ is not in the subtree rooted at $s$, then $\mathsf{Des}(s)\cap\mathsf{Des}(s^{\prime})=\emptyset$, and so $P_{s}P_{s^{\prime}}^{T}=\mathbf{0}$. If $s$ is in the subtree rooted at $s^{\prime}$, then $s$ is in the subtree rooted at some $\widehat{s}\in\mathsf{Chi}(s^{\prime})$. In particular, $(P_{s}P_{s^{\prime}}^{T})(i,j)=0$ for all $j\neq\widehat{s}$ and $(P_{s}P_{s^{\prime}}^{T})(i,\widehat{s})=|\mathsf{Des}(\widehat{s})|^{-1}$, and thus $(I-\mathbf{1}(\frac{|\mathsf{Des}(j)|}{|\mathsf{Des}(s)|})_{j\in\mathsf{Chi}(s)}^{T})P_{s}P_{s^{\prime}}^{T}=\mathbf{0}$. Similarly for the case when $s^{\prime}$ is in the subtree rooted at $s$. The case when $\overline{d}>1$ follows by applying the above argument coordinate-wise.
For the second claim, consider the case $\overline{d}=1$, denoting
@@
where the norm in the denominator in the definition of $v$ is the $l_{2}$ norm. Then when $\overline{d}=1$, we have (applying Theorem 6 in [31])
@@
The above argument can be applied coordinate-wise, and so extend to $\overline{d}\geq 1$. ∎